40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 24 00:39:47 2017, gautam, Update, LSC, DRMI locking attempt 
    Reply  Fri Aug 25 01:20:52 2017, gautam, Update, LSC, DRMI locking attempt 
       Reply  Wed Aug 30 19:49:33 2017, gautam, Update, LSC, REFL55 demod board debugging REFL55_demod_check.pdf
          Reply  Thu Aug 31 00:52:52 2017, gautam, Update, LSC, REFL55 whitening board debugging REFL55_whtCheck.pdfREFL55_whtChk.tar.gz
             Reply  Thu Aug 31 03:31:15 2017, gautam, Update, LSC, DRMI re-locked! DRMI_1f.pngDRMI_relocked.pdf
                Reply  Mon Sep 4 16:30:06 2017, gautam, Update, LSC, Oplev loop tweaking 
                   Reply  Mon Sep 4 18:18:29 2017, rana, Update, LSC, dewhite switching: FOTON settings 
                   Reply  Tue Sep 5 02:07:49 2017, gautam, Update, LSC, Low Noise DRMI attempt MICH_err_comp.pdfdeWhitenedCoil.pdf
                      Reply  Wed Sep 6 23:06:30 2017, gautam, Update, LSC, Coil de-whitening switching investigation BScoils.pdfITMcoils.pdf
                         Reply  Fri Sep 15 16:00:33 2017, gautam, Update, LSC, Sensing measurement DRMI1f_Sep5.pdf
                         Reply  Fri Sep 15 17:08:58 2017, gautam, Update, LSC, Coil de-whitening switching investigation DRMI_IntensityNoise.pdf
                            Reply  Sat Sep 16 10:56:19 2017, rana, Update, LSC, Coil de-whitening switching investigation 
                               Reply  Fri Sep 22 18:12:27 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                  Reply  Wed Sep 27 22:25:21 2017, gautam, Update, LSC, DAC noise measurement (again) coilNoises.pdf
                                     Reply  Thu Sep 28 06:35:44 2017, Chris, Update, LSC, DAC noise measurement (again) 
                                        Reply  Thu Sep 28 23:50:04 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                           Reply  Fri Sep 29 18:36:25 2017, gautam, Update, LSC, DAC noise measurement (again) 
                               Reply  Tue Oct 3 01:32:39 2017, gautam, Update, LSC, Laser intensity noise coupling to MICH (simulated) MICH_intensityNoiseCoupling.pdfMICH_intensityNoiseASD.pdf
Message ID: 13300     Entry time: Wed Sep 6 23:06:30 2017     In reply to: 13291     Reply to this: 13313   13314
Author: gautam 
Type: Update 
Category: LSC 
Subject: Coil de-whitening switching investigation 

Summary:

Rana suggested checking if the coil de-whitening switching is actually happening in the analog path. I repeated the test detailed hereAttachments #1 and #2 suggest that all the coils for the BS and ITMs are indeed switching yes.

Details:

  • The motivation behind this test was the following - the analog path switching is done by applying some logic voltage to a switch, but if this voltage is common among many switches, the hypothesis was that perhaps individual switches were not getting the required voltage to engage the switching.
  • This time FM9 (simulated de-whitening) and FM10 (inverse de-whitening) in the coil output filter modules turned off, so as to maintain a flat TF in the digital domain, but engage the de-whitened analog path (turning off FM9 is supposed to do this).
  • There is poor coherence in the measurement above 40Hz so the data there should be neglected. It is hard to get a good measurement at higher frequencies because of the pendulum TF + heavy low pass filtering from the analog de-whitening path.
  • But between 10-40Hz, we already see the analog de-whitening TF in the measurement.
  • For comparison, I have plotted the measured pendulum TFs for one of the coils from an earlier test (all the coils were roughly at the same level).

So it would seem that there is some other noise which has a 1/f^2 shape and is at the same level we expected the DAC noise to be at. Rana suggested checking coherence with MC transmission to see if this could be laser intensity noise.

I also want to re-do the actuator calibrations for the vertex optics again before re-posting the revised noise budget.

Attachment 1: BScoils.pdf  209 kB  | Hide | Hide all
BScoils.pdf
Attachment 2: ITMcoils.pdf  181 kB  | Hide | Hide all
ITMcoils.pdf
ELOG V3.1.3-